3.55 grams of O2 is enclosed in a 1.5-L container. How many additional grams of O2 should be added to the container to increase the volume to 3.0 L at the same temperature and pressure?
1
Expert's answer
2020-05-11T14:23:46-0400

Volume of gas "(V)\propto" Number of moles of gas "(n) \propto" Mass of gas enclosed(M)

So,"\frac{V_1}{V_2}=\frac{M_1}{M_2}"

"V_1=1.5L,M_1=3.55\ g;V_2=3\ L,M_2=?"

"\frac{1.5}{3}=\frac{3.55}{M_2}" "\implies M_2=3.55\times 2=7.10\ g"

Additional gas to be added"=7.10\ g-3.55\ g=3.55\ g"